find the arc length of f over [2,6] for f(x) =4/3x^3/2


1
Expert's answer
2021-01-10T14:41:06-0500

We have function "f(x) = 4\/3 *" x3/2.and interval [2;6]. We can use one formula for finding arc length: "l = \int_{a}^{b} \sqrt{1+(f'(x))^2} \,dx"

where a, b are ends of the interval, f'(x) is derivative of our function.

So let's find f'(x):

f'(x) = 4/3*3/2*x1/2= 2*x1/2="2*\sqrt{x}"

And find our integral:

"l = \int_{2}^{6} \sqrt{1+(2*\sqrt{x} )^2} \,dx = \int_{2}^{6} \sqrt{1+4x} \,dx="

"= 1\/4*2\/3*(1+4x)^{3\/2} |_{2}^{6}= 1\/6*(1+4x)^{3\/2} |_{2}^{6} ="

"= 1\/6*(25^{3\/2}-9^{3\/2})=1\/6*(125-27)= 98\/6 = 49\/3"

Finally, we have our answer: 49/3 is arc length of f(x) over [2;6], where f(x) =4/3*x3/2